OpenStudy (anonymous):

Alright, here is the Fooool's problem of the day: If for some natural numbers \( A\) and \( B\), \( B^2 +BA+1\) divides \(A^2 +AB +1\), then \( A − B = k\), where \( k \) is an integer. How many such \( k\) exist?

OpenStudy (anonymous):

Am i? To me I look as red as before.

OpenStudy (anonymous):

Duh, I just realized that this is the easiest one of all I posted.

OpenStudy (anonymous):

Sorry not the correct answer. PS: Does programming fails after all? :P

OpenStudy (anonymous):

I am nowhere near the solution.

OpenStudy (anonymous):

And, I am sure this isn't the easiest of all foolformath's questions. or maybe, I am missing something.

OpenStudy (anonymous):

If it helps, I am getting a crazy expression after some substitution and algebra, which can be incorrect too. \[1 + \frac{2kB + k^2}{2B^2 + Bk +1}=n, n \in \mathbb{R^{+}}\]

OpenStudy (mr.math):

Ishaan forced me to sign in to say don't give hints just yet. I'm getting the answer :P

OpenStudy (anonymous):

-sigh- I wish I was a mathematical Genius, do all kinds of calculation and algebra in my head. Sorry lol :-D

OpenStudy (anonymous):

Ishaan, this is not typical algebra.

OpenStudy (mr.math):

I think A has to be equal to B, hence there's only one such k and k=0.

OpenStudy (anonymous):

But what about others? :-D btw I think your solution is right.

OpenStudy (mr.math):

I do have a proof by the way (or so I think). :D

OpenStudy (anonymous):

Yes, that's right. Let us see the proof Mr.Math.

OpenStudy (mr.math):

Sure! :)

OpenStudy (mr.math):

Alright! First, someone can easily see that \(A^2+AB+1>0\) and \(B^2+AB+1>0\), for \(A, B\in \mathbb{N}\). So it follows that \(A^2+AB+1\ge B^2+AB+1 \implies A\ge B\), and we can write: \[cA^2+cAB+c=B^2+AB+1, \text{ where c is an integer}\ge 1. \] From that last equation we can get: \[(c-1)A^2+A^2-B^2+(c-1)AB+c-1=0\] \[\implies (c-1)(A^2+AB+1)+(A-B)(A+B)=0,\] \[ \text{ but since A-B=k, we can write}:\] \[k=\frac{(1-c)(A^2+AB+1)}{A+B}.\] That's (A+B)|(A^2+AB+1) or (A+B)|(1-c): Since \(\frac{A^2+AB+1}{A+B}=\frac{1}{A+B}+1\), then there's no natural numbers A,B such that (A+B)|(A^2+AB+1). So, it has to be the case that (A+B)|(1-c). The obvious solution here is that 1-c=0 or c=1. Hmm, I think there's still something missing here.

OpenStudy (mr.math):

Yeah, I got what's missing. We have shown above that A>=B, that k>=0, which implies that \[1-c\ge 0 \implies c\le 1\] Thus c=1 is the only solution, which gives k=0.

OpenStudy (mr.math):

Is the proof clear?

OpenStudy (anonymous):

You lost me after the first line itself :P

OpenStudy (mr.math):

A and B are natural then A^2+AB>=1, right?

OpenStudy (mr.math):

In other words both the top and bottom are positive, so we need not to worry about negative divisors (i.e the top has to be greater than or equal to the bottom).

OpenStudy (mr.math):

And it follows that A>=B.

OpenStudy (anonymous):

Yes yes, I mean't this was not the elementary proof that I was talking about :)

OpenStudy (mr.math):

Oh, I see. I know I always complicate things :P

OpenStudy (anonymous):

It's so simple, If A divides B then A divides B-A, and then ... ;)

OpenStudy (mr.math):

I like my proof! :P

OpenStudy (mr.math):

Ishaan likes it too (I think).

OpenStudy (anonymous):

If I post mine yours will be rendered too complicated :P

OpenStudy (mr.math):

Maybe! :D

OpenStudy (mr.math):

Ahh, I have to get back to study group theory! :(

OpenStudy (mr.math):

Honestly, I didn't know (or consider) that if A|B then A|(B-A).

OpenStudy (anonymous):

That's Number theory 101 for any Mathelete ;)

OpenStudy (mr.math):

Yeah, when I think about it. Its proof is trivial too.